# Duplex-GS: Proxy-Guided Weighted Blending for Real-Time Order-Independent Gaussian Splatting

### v1

@@ Abstract
Recent advances in 3D Gaussian Splatting (3DGS) have demonstrated remarkable rendering fidelity and efficiency. However, these methods still rely on computationally expensive sequential $\alpha$-blending operations, resulting in significant overhead—particularly on resource-constrained platforms. In this paper, we propose Duplex-GS, a dual-hierarchy framework that integrates proxy Gaussian representations with order-independent rendering techniques to achieve photorealistic results while sustaining real-time performance. To mitigate the overhead caused by view-adaptive radix sort, we introduce cell proxies for local Gaussians management and propose cell search rasterization for further acceleration. By seamlessly combining our framework with Order-Independent Transparency (OIT), we develop a physically inspired weighted sum rendering technique that simultaneously eliminates “popping” and “transparency” artifacts, yielding substantial improvements in both accuracy and efficiency. Extensive experiments on a variety of real-world datasets demonstrate the robustness of our method across diverse scenarios, including multi-scale training views and large-scale environments. Our results validate the advantages of the OIT rendering paradigm in Gaussian Splatting, achieving high-quality rendering with an impressive $1.5\times$$4\times$ speedup over existing OIT based Gaussian Splatting approaches and $52.2\%\,$$86.9\%$ reduction of the radix sort overhead without quality degradation.
Index Terms:
Novel View Synthesis, Gaussian Splatting, Order-Independent Transparency, Real-Time Rendering
@@ I Introduction
Novel view synthesis (NVS) and real-time rendering have attracted significant attention in both academia and industry, with broad applications in video games, virtual reality (VR) [1], and 3D scene reconstruction [2, 3, 4, 5]. Inspired by the groundbreaking Neural Radiance Field (NeRF) [6], learning based methods have substantially advanced the realism of digital content, narrowing the gap between the virtual and real worlds. However, these methods often entail considerable computational overhead. The emergence of 3D Gaussian Splatting (3DGS) [7, 8] marks another milestone, enabling photorealistic and real-time rendering of radiance fields via rasterization—a paradigm akin to the graphics pipeline in modern graphics processing units (GPUs).
@@
Despite the impressive reconstruction and rendering capabilities of 3DGS, a primary limitation is the need for an ever-increasing number of Gaussian primitives to fit all training views. This results in substantial and irregular memory access during rendering, as costly view-dependent sorting operations over long sequences impose considerable computational and memory overhead [9, 10]. Moreover, the sorting process can introduce visible “popping” artifacts [11], as illustrated in Fig. 1. Collectively, these limitations hinder the scalability of 3DGS for large-scale applications and restrict its deployment on resource-constrained platforms such as mobile phones and VR devices.
@@
More recently, sort-free Gaussian Splatting [12] has incorporated Order-Independent Transparency (OIT) [13, 14] into the Gaussian Splatting framework, eliminating the need for sorting in the rasterization pipeline by employing Weighted Sum Rendering (WSR) [15]. This effectively eliminates “popping” artifacts and has been successfully demonstrated on mobile devices [12]. However, as illustrated in Fig. 1, the absence of early termination when encountering opaque objects leads to “transparency” artifacts and increased rendering time, particularly in large-scale scenes, due to the accumulation of redundant Gaussian primitives in final blending.
Neural Gaussian framework [16, 17] is another line of work to enhance storage efficiency for 3DGS, which transitions from explicit 3D Gaussian primitives to a hybrid representation where clusters of Gaussians are encoded via “anchors” and associated with a shared MLP. However, this paradigm adheres to the conventional rendering pipeline and still operates on sets of 3D Gaussians decoded on-the-fly by the MLP. As a result, sorting the lengthy sequence of Gaussians remains mandatory for such Tile based Rendering (TBR), presenting substantial challenges for devices with limited resources [10]. Furthermore, since the spatial positions of neural Gaussians are not explicitly constrained by the anchor points, the generated Gaussians may deviate significantly from their associated anchors—as illustrated in Fig. 2—thereby complicating anchor management.
Figure 1: Motivation of Duplex-GS. The rendering process for 3DGS rely on sorting for long tile-Gaussian sequence, introducing popping artifacts. The sort-free LC-WSR uses weighted sum rendering method without the need of consuming sorting stage, introducing transparency artifacts due to the lack of early termination when encounter opaque object.
Figure 2: Overview of the proposed Duplex-GS. The dual-hierarchy structure introduce cell as the proxy of local similar Gaussians. (a) Compared with proxy anchor, the proposed proxy cell is an ellipsoid with specific geometry. (b) Combing OIT rendering with 3DGS models, our method successfully remove the “popping” and “transparency” artifacts simultaneously. (c) By introducing cell search rasterization (Sec. IV) and physical weighted sum rendering (Sec. V), we achieve significant improvement over OIT Gaussian Splatting for both accuracy and efficiency.
In this paper, we present Duplex-GS, which models 3D scenes with WSR rendering paradigm and dual-layered structure following the neural Gaussian framework. Firstly, scenes are encoded as explicit proxies that are easily stored and manipulated in the first layer. As the concept anchor degenerates the 3D Gaussian to point cloud, we introduce “cell” as the replacement, which is an ellipsoid with concrete physical geometry initialized at the Structure-from-Motion (SfM) [18] points. By constraining the emitted neural Gaussians located inside the cell region, the rasterization can be performed in cell level. Secondly, primitives are decoded as other forms, which is suitable for high-fidelity rendering in the second layer. Encouraged by [12], we introduce the physical inspired WSR with early termination where the specific order of Gaussian depth is not required, which replaces the non-commutative blending operations with parallelizable weighted sum operation. As the number of primitives in the first layer is extensively fewer than that of Gaussians, the proposed cell search rasterization significantly reduce the memory and computational overhead introduced by radix sort. On the one hand, rasterization and lightweight sorting in the first layer removes the “transparency” artifact. On the other hand, the WSR kernel removes the “popping” artifact introduce by vanilla $\alpha$-blending.
In summary, the contributions of this work are as follow:
  • We introduce a dual-layered Gaussian hierarchy, in which the first layer consists of proxies called cells and the second layer comprises 3D Gaussians. Our approach significantly alleviates the need for global sorting while preserving the parallelizability of WSR, thereby achieving artifact-free, real-time 3D reconstruction.
@@
  • Cell rasterization in layer I: We introduce cells—ellipsoids with adaptive geometric properties—to organize emitted Gaussians. Based on this structure, we propose a novel cell rasterization strategy that dramatically reduces both memory and computational overhead associated with radix sort.
@@
  • Physical inspired WSR blending in layer II: We retrieve physical constraints of Gaussian parameters for WSR method to blend Gaussians without the need for global sorting, inherently avoiding the “popping” artifact. Building on this, we introduce cell-level early termination to further accelerate rendering and effectively mitigate the “transparency” artifact encountered in the original WSR approach.
@@
  • We implement WSR based Gaussian Splatting on CUDA-enabled platforms. Experimental results demonstrate that our method delivers comparable rendering speed and visual fidelity to existing approaches on commercial high-end GPUs.
@@ II-A Neural Representation for 3D Scene
Traditional scene reconstruction methods, such as SfM [18, 19] and Multi-View Stereo (MVS) [20, 21], often struggle to recover texture-less regions or handle challenging lighting variations. The advent of Neural Radiance Fields (NeRF) [6] has demonstrated the impressive capability of learning based models to achieve photorealistic reconstructions from sparse 2D images. Building on this, numerous efficient NeRF variants have been developed [22, 23, 24], making real-time rendering increasingly feasible on edge devices [25]. A significant breakthrough in the field is 3D Gaussian Splatting (3DGS) [7], which leverages explicit anisotropic 3D Gaussians for scene reconstruction and substantially accelerates rendering. Nevertheless, the substantial computational and memory demands of 3DGS continue to limit its applicability on resource-constrained platforms. Notably, a trend is witnessed that as scene representations become more explicit, rendering speed improves at the cost of increased memory requirements. And hybrid variants have emerged as a promising compromise between rendering efficiency and memory usage [22, 26, 16]. For instance, Instant-NGP [22] employs an explicit learnable hash table that maps 3D space to a multi-resolution grid, thereby alleviating the computational burden on the MLP. Scaffold-GS [16] introduces a hierarchical structure that encodes neural Gaussians using anchor points, achieving comparable rendering quality and speed with a significantly reduced model size. These advances enable fast, high-resolution rendering without compromising visual fidelity. However, these approaches still rely on the conventional blending pipeline requiring the Gaussians to be rendered in sorted depth order, typically achieved with tile based sorting algorithms. This sorting step not only complicates practical implementation but also introduces visual artifacts such as sudden changes in appearance, commonly referred to as “popping” artifacts [11].
@@ II-B Efficient Neural Rendering
Unlike NeRF whose constraint is the high computational demand, the bottlenecks of 3DGS practical deployment are the extensive number of Gaussian primitives and the frequent memory access requirements [10, 9]. Various works aim to improve the efficiency of 3DGS by reducing the model size or the number of Gaussians involved in the sorting based blending operations [17, 27, 28, 29]. Encouraged by the OIT [13, 14], a widely deployed rendering paradigm in modern GPU to blend non-opaque media, sort-free Gaussian Splatting [12] introduce a novel blending paradigm WSR which eliminates the need of sorting operation in 3DGS, which removes the “popping” artifact as well as achieves faster rendering speed on mobile devices with small scale scenes. However, the lack of early termination results in the “transparency” artifacts as well as introduces extra computation when encounter opaque objects.
@@ II-C Order Independent Transparency
Order-Independent Transparency (OIT) [13, 14] is a foundational graphics technique integrated into modern GPUs to enable the correct compositing of transparent fragments without requiring explicit depth sorting prior to rasterization. Early approaches, such as Depth Peeling [30], A-buffer [31], and Multi-Layer Alpha Blending (MLAB) [32], successfully achieve OIT but often incur significant time or memory overhead. Weighted Blended OIT [15] offers a more efficient alternative by approximating the integration of transparency in a single render pass, storing weighted sums of color and opacity with carefully designed blending weights. And this method has been applied to 3D Gaussian Splatting (3DGS) rendering [12]. However, the absence of early termination in [12] when encountering opaque regions can degrade both visual quality and rendering performance. In contrast, the proposed Duplex-GS framework integrates the core idea of early termination from depth peeling—halting compositing upon encountering fully opaque surfaces—with Weighted Blended OIT. As a result, Duplex-GS attains rendering quality and speed on par with state-of-the-art (SOTA) 3DGS models on high-end GPUs, while exhibiting strong potential for deployment on resource-constrained edge devices by alleviating the huge overhead of the view-adaptive sorting stage.
Figure 3: Illustration of the proposed Duplex-GS pipeline. A dual-layered structure is designed to organize cells and Gaussians. In Layer I, proxy cells—represented as ellipsoids with specific geometric properties—are designed to aggregate Gaussians, enabling rasterization at the cell level and thereby alleviating the overhead associated with view-adaptive global sorting. Leveraging shared MLPs, the cell primitives are subsequently decoded into Gaussian primitives. In Layer II, the decoded Gaussians are blended using weighted sum rendering. Additionally, early termination technique is implemented—guided by the sorting results of the sparse cells from Layer I—which facilitates real-time, photorealistic rendering.
III Preliminaries
III-A Gaussian Splatting
3D Gaussian Splatting [7] models scenes using anisotropic 3D Gaussian primitives and renders images through a rasterization based paradigm. Each 3D Gaussian $G(\bm{x})$ is represented by a center coordinate $\bm{\mu}\in\mathbb{R}^{3}$ and a covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$:
$G(\bm{x})=e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})},$ (1)
where $\bm{x}\in\mathbb{R}^{3}$ is the queried position of the Gaussian kernel, and the covariance $\bm{\Sigma}$ is parameterized by scales $\bm{s}\in\mathbb{R}^{3}$ and quaternions $\bm{q}\in\mathbb{R}^{4}$. In line with volumetric rendering conventions, each Gaussian is further associated with an opacity $\sigma\in\mathbb{R}$ and a view-dependent color $\bm{c}\in\mathbb{R}^{3}$, which is commonly encoded using Spherical Harmonics (SH). Rendering proceeds via a tile based Gaussian rasterization technique, which employs efficient radix sort to enable $\alpha$-blending
$\bm{C}(\bm{x}^{\prime})=\sum_{i\in N}T_{i}^{\text{3DGS}}\bm{c}_{i}\alpha_{i},\quad\alpha_{i}=\sigma_{i}G^{\prime}_{i}(\bm{x}^{\prime}),$ (2)
where $\bm{x}^{\prime}$ denotes the 2D coordinates of the queried pixel, $G^{\prime}_{i}(\bm{x}^{\prime})$ is the projected 2D gaussian kernel, and $T_{i}^{\text{3DGS}}$ denotes the transmittance
$T_{i}^{\text{3DGS}}=\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (3)
III-B Scaffold-GS
@@
Scaffold-GS introduces anchors as local encoding proxies, which emit neural Gaussians decoded from latent features to effectively represent local scene structure. Specifically, each anchor emits a predefined set of $K$ neural gaussians, whose centers are given by:
@@
where $F_{\sigma}$ is the shared opacity decoder. Similarly, the other attributes of each neural Gaussians, including color $\bm{c}_{n,k}$, scales $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$, are predicted by the separate MLPs $F_{c}$, $F_{s}$, $F_{q}$, respectively.
III-C Weighted Sum Rendering
Inspired by Order Independent Transparency (OIT)—a widely adopted technique for rendering transparent media in traditional graphics pipelines—sort-free Gaussian Splatting integrates this concept into 3DGS, producing the final image as

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) achieves photorealistic rendering but incurs substantial overhead due to the global sorting required for $\alpha$-blending, which results in noticeable “popping” artifacts and hinders deployment on edge devices. While sort-free Order-Independent Transparency (OIT) methods circumvent sorting, they introduce “transparency” artifacts and suffer from inefficiency due to the absence of physical constraints. To address these limitations, we present Duplex-GS, a dual-hierarchy framework that leverages proxy-guided spatial organization and a novel hybrid renderer that combines $\alpha$-blending with reformulated Weighted-Sum Rendering (WSR). We introduce explicit ellipsoidal cell proxies to encapsulate local Gaussians, which enables efficient proxy-level rasterization. This strategy drastically reduces the overhead associated with global sorting. Furthermore, we propose a physically grounded WSR scheme with cell-level early termination, which retrieves the physical constraints absent in prior OIT-based 3DGS methods, effectively eliminating both popping and transparency artifacts. Extensive experiments on diverse real-world datasets demonstrate the robustness of Duplex-GS across scenarios spanning multi-scale training views and large-scale environments. Quantitatively, our method delivers high-fidelity real-time rendering, outperforming prior OIT-based 3DGS methods by $1.5\times$– $4\times$ in speed, while reducing radix-sort cost by $29.8\%\,$– $86.9\%$ compared with conventional $\alpha$-blending without compromising visual quality.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
Novel view synthesis (NVS) and real-time rendering have attracted significant attention in both academia and industry, with broad applications in video games, virtual reality (VR) [1], and 3D scene reconstruction [2, 3, 4, 5, 6, 7, 8, 9]. Inspired by the groundbreaking Neural Radiance Field (NeRF) [10], learning-based methods have substantially advanced the realism of digital content, narrowing the gap between the virtual and real worlds. However, these methods often entail considerable computational overhead. The emergence of 3D Gaussian Splatting (3DGS) [11, 12] marks another milestone, enabling photorealistic and real-time rendering of radiance fields via rasterization, a paradigm akin to the graphics pipeline in modern graphics processing units (GPUs).
@@
Despite the impressive reconstruction and rendering capabilities of 3DGS, a primary limitation is its need for an ever-increasing number of Gaussian primitives to fit all training views. This results in substantial and irregular memory access during rendering, as costly view-dependent sorting operations over long sequences impose considerable computational and memory overhead [13, 14]. Moreover, the sorting process can introduce visible “popping” artifacts [15], as illustrated in Fig. 1. Collectively, these limitations hinder the scalability of 3DGS for large-scale applications and restrict its deployment on resource-constrained platforms such as mobile phones and VR devices.
@@
More recently, sort-free Gaussian Splatting [16] has incorporated Order-Independent Transparency (OIT) [17, 18] into the Gaussian Splatting framework, eliminating the need for sorting in the rasterization pipeline by employing Weighted Sum Rendering (WSR) [19]. This effectively eliminates popping artifacts and has been successfully demonstrated on mobile devices [16]. However, as illustrated in Fig. 1, the absence of physical constraints leads to “transparency” artifacts and increased rendering time, particularly in large-scale scenes, due to the accumulation of redundant Gaussian primitives and lack of early termination when encountering opaque objects in the blending phase.
Neural Gaussian frameworks [20, 21] enhance storage efficiency by using “anchor” points to encode local Gaussians via shared MLPs. However, they still require sorting the decoded Gaussians for $\alpha$-blending, which is costly on resource-constrained devices [14]. Moreover, the lack of explicit spatial constraints allows Gaussians to deviate from their anchors (Fig. 2), complicating the representation.
Figure 1: Motivation of Duplex-GS. (Top) 3DGSs are rendered via $\alpha$-blending for a long tile-Gaussian sequence. However, sudden changes in viewing angle can cause abrupt changes in the blending order and the corresponding weights, resulting in noticeable popping artifacts. (Bottom) The sort-free LC-WSR method eliminates the sorting stage through WSR paradigm. However, the absence of physical constraints leads to transparency artifacts and efficiency degradation, as blending continues unnecessarily for opaque surfaces.
Figure 2: Overview of the proposed Duplex-GS. Our approach is built on a dual-hierarchy structure that introduces ellipsoidal cells as proxies for local Gaussians. (a) Explicit Cell Proxies. Unlike abstract anchor points, our geometrically defined proxy cells are directly utilized for rasterization, guiding the subsequent Gaussian blending process. (b) Hybrid Rendering Paradigm. By synergistically combining $\alpha$-blending with physically grounded WSR, our method simultaneously eliminates both popping and transparency artifacts. (c) The integration of cell rasterization (Sec. IV) and physically grounded WSR (Sec. V) achieves significant improvements over prior OIT-based Gaussian splatting methods in both rendering accuracy and computational efficiency.
This paper presents Duplex-GS, a novel framework for modeling 3D scenes using explicit proxies with defined visible regions and a tailored hybrid rendering paradigm. Our approach integrates the benefits of $\alpha$-blending and WSR to eliminate artifacts while maintaining real-time performance. First, we encode scenes using ellipsoidal proxies, termed “cells”, which replace the anchors used in the original neural Gaussian framework. By constraining neural Gaussians to reside within their respective cells, rasterization is performed at the cell level to guide subsequent Gaussian blending. Second, activated cells are decoded into Gaussians via shared MLPs. By restoring physical constraints absent in previous OIT-based 3DGS [16], we introduce a physically grounded WSR strategy that enables early termination and replaces non-commutative blending operations with a parallelizable weighted sum.
This hybrid design offers two key advantages. First, the significantly smaller population of cell proxies compared to Gaussians enables our cell-level rasterization strategy to drastically reduce the memory and computational overhead of global radix sorting. Second, we restore physical constraints absent in prior OIT-based 3DGS method, which facilitates cell-level early termination and removes both popping and transparency artifacts simultaneously.
In summary, the contributions of this work are as follows:
  • We propose a dual-layered Gaussian hierarchy comprising ellipsoidal proxy cells and 3D Gaussians. This structure enables artifact-free, real-time 3D reconstruction while maintaining the parallelizability of WSR.
@@
  • Cell rasterization (layer I): We introduce ellipsoidal cells with adaptive geometric properties to spatially organize neural Gaussians. A novel cell-level rasterization strategy is proposed, which significantly reduces the memory and computational overhead associated with global radix sorting.
@@
  • Physically grounded WSR blending (layer II): We incorporate physical constraints into the WSR method to blend local Gaussians, inherently avoiding popping and transparency artifacts. Building on this, we introduce a cell-level early termination scheme that further accelerates rendering.
@@
  • We provide a high-performance CUDA implementation of our WSR-based Gaussian splatting framework. Experimental results demonstrate that our method achieves competitive rendering speed and superior visual fidelity on consumer-grade hardware.
@@ II-A Neural Representation for 3D Scene
Traditional scene reconstruction methods like SfM [22, 23] and MVS [24, 25] struggle with texture-less regions and lighting variations. NeRF [10] enabled photorealistic reconstruction from sparse images, with subsequent efficient variants [26, 27, 28] making real-time rendering feasible on edge devices [29, 30]. 3DGS [11] marked a breakthrough by using explicit anisotropic Gaussians for fast rendering, though its computational and memory demands limit its deployment on resource-constrained platforms. A clear trend shows that more explicit representations improve speed at the cost of memory, leading to hybrid approaches [26, 31, 32, 20, 33, 34, 35] that balance this trade-off. For example, Instant-NGP [26] uses a multi-resolution hash table to reduce MLP burden, Scaffold-GS [20] encodes neural Gaussians via anchors to reduce model size while maintaining quality. However, these methods still rely on the conventional $\alpha$-blending pipeline that requires sorting Gaussians by depth. This step complicates implementation and introduces visual popping artifacts [15].
@@ II-B Efficient Neural Rendering
While NeRF is primarily constrained by high computational demands, the practical deployment of 3DGS is bottlenecked by its extensive number of Gaussian primitives and frequent memory access requirements [14, 13]. Existing efforts to optimize the original $\alpha$-blending pipeline in 3DGS broadly fall into three categories: 1) Applying model compression techniques, such as Level-of-Detail (LOD) strategy [21, 36, 37, 38], encoding schemes [20, 39, 40], as well as pruning and quantization [41, 42, 43]. 2) Optimizing critical operators through CUDA-based implementations [44, 45, 46, 47, 48] or dedicated hardware acceleration units [49]. 3) Replacing Gaussian kernels with more efficient representation such as linear kernals [50] and half-Gaussian kernel [51].
A few works have attempted to revise the core rendering paradigm itself by exploring efficient alternatives to $\alpha$-blending. StochasticSplats [52] introduces a stochastic transparency paradigm which enables efficient rendering using OpenGL shaders at the risk of controllable quality degradation. Sort-free Gaussian Splatting [16] introduces a Weighted Sum Rendering (WSR) pipeline, which eliminates popping artifacts and achieves faster rendering on mobile devices for small-scale scenes. Nevertheless, its lack of physical constraints leads to transparency artifacts and introduces unnecessary computation when rendering opaque surfaces.
@@ II-C Order Independent Transparency
OIT [17, 18] is a foundational graphics technique integrated into modern GPUs to enable the correct compositing of transparent fragments without requiring explicit depth sorting prior to rasterization. Early approaches, such as Depth Peeling [53], A-buffer [54], and Multi-Layer Alpha Blending (MLAB) [55], successfully achieve OIT but often incur significant time or memory overhead. Weighted Blended OIT [19] offers a more efficient alternative by approximating the integration of transparency in a single render pass, storing weighted sums of color and opacity with carefully designed blending weights. And this method has been applied to 3DGS rendering [16]. However, the absence of physical constraints in[16] degrades both visual quality and rendering efficiency. In contrast, the proposed Duplex-GS framework restores the physically grounded transmittance in proxy-level, enabling early termination to halt compositing upon encountering fully opaque surfaces with weighted blended OIT. As a result, Duplex-GS attains rendering quality and speed on par with state-of-the-art (SOTA) 3DGS models on high-end GPUs, while exhibiting strong potential for deployment on resource-constrained edge devices by alleviating the huge overhead of the view-adaptive sorting stage.
Figure 3: Illustration of the proposed Duplex-GS pipeline. (Layer I) Ellipsoidal cells act as proxies to spatially organize local Gaussians. This structure enables efficient cell-level rasterization, significantly reducing the overhead of view-adaptive global sorting. (Layer II) Intersected cells are decoded into Gaussians and blended with the proposed physically grounded WSR. The final color is computed by first determining cell-level weights via $\alpha$-blending, then aggregating the Gaussians inside each cell in an order-independent WSR manner.
III Preliminaries and Overview
III-A Preliminaries
III-A1 Neural Gaussian
@@
3DGS [11] models scenes using anisotropic 3D Gaussian primitives and renders images through a rasterization based paradigm. Building upon 3DGS, Scaffold-GS [20] introduces anchors as local encoding proxies, which emit neural Gaussians decoded from latent features to effectively represent local scene structure. Specifically, each anchor emits a predefined set of $K$ neural gaussians, whose centers are given by:
@@
where $F_{\sigma}$ is the shared opacity decoder where the output is activated by Tanh. Similarly, the other attributes of each neural Gaussians, including color $\bm{c}_{n,k}$, scales $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$, are predicted by separate MLPs $F_{c}$, $F_{s}$, $F_{q}$, respectively. Consequently, rendering proceeds via a tile based Gaussian rasterization technique, which employs radix sort to blend $M$ ordered points
$\bm{C}(\bm{x}^{\prime})=\sum_{i\in M}T_{i}^{\text{$\alpha$-blending}}\bm{c}_{i}\alpha_{i},\quad\alpha_{i}=\sigma_{i}G^{\prime}_{i}(\bm{x}^{\prime}),$ (3)
where $\bm{x}^{\prime}$ denotes the 2D coordinates of the queried pixel, $G^{\prime}_{i}(\bm{x}^{\prime})$ is the projected 2D gaussian kernel, and $T_{i}^{\text{$\alpha$-blending}}$ denotes the transmittance
$T_{i}^{\text{$\alpha$-blending}}=\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (4)
III-A2 Weighted Sum Rendering
Inspired by OIT, a widely adopted technique for rendering non-opaque media in traditional graphics pipelines, sort-free Gaussian Splatting integrates this concept into 3DGS, producing the final image as